5x/7 + 2= x/2 -4
This is 5x over 7 as a fraction plus 2= x over 2 as a fraction minus 4.
Solve for x

Answer by Juhi(44)

5x/7 - x/2 = -4 - 2
3x/14 = -6
x = 14/3 * 6
= 14 * -2
= -28